OpenStudy (anonymous):

Simplify the expression.. 64 1/3 @Opcode

OpenStudy (opcode):

Um, Is it: \[64^{1/3}\] Or \[\dfrac{64^1}{3}\] Do you have answer choices?

OpenStudy (anonymous):

The first one, sorry.

OpenStudy (anonymous):

Also answer choices? No. It's just a type the solution thing.. I'll take a screenshot, hold on.

OpenStudy (anonymous):

it is the first on k

OpenStudy (anonymous):

@Opcode Do you know how to help me with this mate? :)

OpenStudy (anonymous):

yea

OpenStudy (opcode):

Basically, you root the base (in this case 64) by the bottom number of the fraction (here it is 3), and then put that to the power of the top number (1). \[64^3 = 262144\] \[4th \sqrt{262144} = 22.62\] I cannot tell if I did that right or wrong sorry. The way to get the fourth root is by square rooting twice.

OpenStudy (anonymous):

It was wrong ._. yeah.. I don't know how they want me to do that either. It's not too helpful when all they say is "Simplify the expression" .. It'd be nice if they said "Simplify the expression by ____" .. lol. But thanks anyways bro.

OpenStudy (opcode):

Sorry, mate. Well I don't see any other way to do that question :-/. Maybe that wanted the answer rounded? Since it is 22.62 the six rounds it to 23?

OpenStudy (anonymous):

Possibly.

OpenStudy (opcode):

Well this sucks, I'll ask my teacher tomorrow on how to do stuff like that.